29- Auerbach Inc. issued 6% bonds on October 1, 2021. The bonds have a maturity date of September 30, 2031 and a face value of $440 million. The bonds pay interest each March 31 and September 30, beginning March 31, 2022. The effective interest rate established by the market was 8%.

Assuming that Auerbach issued the bonds for $380,204,000, what would the company report for its net bond liability balance after its first interest payment on March 31, 2022?

Multiple Choice

  • $382,212,160.

  • $389,580,360.

  • $382,960,320.

  • $385,167,000.
Answer #1
Discount on issue of Bonds =$440,000,000 - $380,204,000 =$59,796,000
Interest paid on Bond issued =$440,000,000*6%*6/12 =$13,200,000
Interest expenses for 6 month ending March 31,2022 =$380,204,000*8%*6/12 =$15,208,160
Discount amortized for 6 month ending March 31,2022 =$15,208,160 - $13,200,000 =$2,008,160
Net Bond liability as on March 31,2022 =$380,204,000 + 2,008,160 =$382,212,160
So Option A is the answer